TWO CARS LEAVE AN INTERSECTION AT THE SAME TIME. ONE TRAVELS NORTH AND THE OTHER HEADS SOUTH TRAVELING 10MPH FASTER. AFTER ONE HOUR, THEY ARE 50 MILES APART. WHAT WERE THEIR SPEEDS?
Answer by mananth(16949)

they are traveling in opposite directions.
let one car travel at xmph
the other travels at x+10 mph
...
Since they are traveling in opposite direction ,colinesr add up the speeds.
Their effective speed = x+1+10 mph= 2x+10 mph.
...
time = distance /speed
time = 1 hour
speed = 2x+10 mph
distance = 50 miles
...
1 = 50/(2x+10)
2x+10 = 50
2x=50-10
2x=40
x= 20 mph
...
x+10 = 30 mph.
...
C H E C K
After 1 hour they will be 20+30 miles apart. = 50 miles.
...
m.ananth@hotmail.ca